OpenStudy (anonymous):

Please help me! Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (-8, 0), (1, 5)

OpenStudy (nubeer):

m= 5-0/1-(-8) m= 5/9 y-0 = 5/9(x-(-8)) y = 5/9 (x+8) 9y = 5x + 40 compare this with Ax+By = C.. you will get values ..

OpenStudy (anonymous):

Equation of Line with one point and slope given is given by : \(y - y_1 = m(x - x_1)\)

OpenStudy (anonymous):

But @nubeer that is what I had as my answer.

OpenStudy (nubeer):

hmm so is it wrong?

OpenStudy (anonymous):

Yes, I think so..

OpenStudy (nubeer):

well I don't think so .. or maybe just rearrage it but that should be the equation of line..

OpenStudy (anonymous):

OK. Thank you!

OpenStudy (anonymous):

hi guys!, m also hvig same problm here @numbeer hw do i rearrage it?

OpenStudy (nubeer):

9y = 5x +40 -5x +9y =40 A = -5 , B = 9 , C = 40

OpenStudy (nubeer):

@mapty

OpenStudy (anonymous):

thnk u nubeer.